# TS-PEFT: Unveiling Token-Level Redundancy in Parameter-Efficient Fine-Tuning

### v2

@@ 3.4 Adam-style Threshold Update and Overall Training Procedure
In practice, the approximate gradient $g_{k}$ for the threshold $\tau$ is highly noisy and varies in scale across training steps: the contributions of tokens change significantly between batches, and the magnitude of $g_{k}$ can fluctuate during optimization. Directly applying plain gradient descent to $\tau$ therefore tends to cause oscillation and slow or unstable convergence. To mitigate this, we update each threshold with an Adam-style rule that combines momentum and adaptive learning rateskinga2015method.
@@
where $\beta_{1}$ controls the contribution of past gradients and $\beta_{2}$ governs the decay rate of squared gradients, while $\hat{m}_{k}$ and $\hat{v}_{k}$ denote the bias-corrected first- and second-moment estimates used in the Adam update. Finally, the threshold is updated as:
@@
where $s$ is the gradient scaling factor, and $\epsilon$ prevents division by zero. In practice, this Adam-style update yields stable convergence of the thresholds under noisy gradients while preserving the desired sparsity level.
Optimization Stability. We empirically observe that the adaptive moment estimation (Eq. (16)-Eq. (20)) is critical for convergence. Ablation studies in Appendix A.2 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks like BoolQ.
@@
Combining the above components, the training process of TS-PEFT can be summarized as follows: in each training step, during the forward pass, we perform gating based on each token’s relative update magnitude $r_{i}$ and the threshold $\tau$ of the corresponding layer, applying updates only at positions that satisfy the condition while keeping the backbone output unchanged at all other positions. Then, standard backpropagation is used to update all PEFT parameters, while the backbone parameters remain frozen throughout. On this basis, we compute $\mu_{i}$ to construct the approximate gradient $g_{k}$ for the threshold, and independently update the threshold $\tau$ of each layer using Adam-style first- and second-moment estimates. The full training procedure is presented in Algorithm 1.
@@ 4.1 Experimental Setup
We evaluate TS-PEFT across three representative scenarios: natural language understanding (NLU), commonsense reasoning (CSR), and visual instruction tuning (VIT). The experimental setup covers the datasets, baseline models, and hyperparameter configurations to ensure reproducibility and comparability of the results.
@@
We select several representative benchmarks in the current PEFT and small-model fine-tuning literature. For CSR, we evaluate on LLaMA-3.1-8B dubey2024llama using PIQA, BoolQ, HellaSwag, WinoGrande, SIQA, OBQA, and ARC-e/ARC-c, covering typical language reasoning scenarios such as physical commonsense, Boolean question answering, narrative completion, and scientific QA liu2024dora. For VIT, we evaluate LLaVA-1.5-7B liu2023visual; liu2024improved on ScienceQA, POPE, MMBench, GQA, VisWiz, VQAv2, and VQAT, which are generative language tasks with open-ended answers conditioned on images and textual instructions, rather than fixed-choice classification. For NLU, we build on DeBERTaV3-base hedebertav3 and use the GLUE benchmark datasets wang2018glue. The GLUE tasks follow the official evaluation protocol: CoLA uses the Matthews correlation coefficient, STS-B uses correlation coefficients, and all other tasks use accuracy.
@@
In our experiments, we select several mainstream parameter-efficient fine-tuning methods as baselines, covering two major categories: low-rank update methods and adapter-based methods. Specifically, LoRA, DoRA, and AdaLoRA are used to examine whether TS-PEFT can consistently provide additional gains under different low-rank update strategies, while Parallel Adapter hu2023llm is included to evaluate the method’s scalability on adapter-style architectures. For each baseline, we keep its original implementation, hyperparameters, and rank configuration unchanged, and apply TS-PEFT on top of it. This ensures that the comparison is not affected by factors unrelated to the baseline itself and that the results accurately reflect the impact of token-level selective updates.
@@
In all experiments, we freeze the backbone model and update only the PEFT parameters and the threshold $\tau$. We use the AdamW optimizer, where the PEFT parameters and $\tau$ share the same base learning rate, while the effective learning rate of $\tau$ is further modulated by the scaling factor s. For text tasks, we follow standard instruction-tuning configurations; for vision–language tasks, we adopt the original training pipeline of the corresponding models. Unless otherwise specified, we set $\alpha_{k}=1$, $\beta_{1}=0.9$, and $\beta_{2}=0.98$. Each main experiment is repeated at least three times with different random seeds, and we report the mean scores across runs. Additional hyperparameter settings are summarized in Appendix A.4.
@@ 4.2 Experimental Results
Commonsense Reasoning. Table 1 reports the results of various PEFT methods and their TS-PEFT variants on the CSR tasks. For a fair comparison, we set the rank of both LoRA and DoRA to 32, and configure AdaLoRA with an initial rank of 64 and a target rank of 32. Under the same rank settings, we observe that TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve performance comparable to or better than their respective baselines on most datasets. This indicates that selectively skipping updates on a subset of tokens does not weaken downstream performance; instead, it can help mitigate the interference caused by redundant updates.
The table also reports the sparsity rate of TS-PEFT (i.e., the proportion of positions in the input sequence that are not modified by the PEFT output, as defined in Eq.(21)). Across different baselines and tasks, this proportion remains roughly within the 40%–60% range, meaning that fewer than half of the tokens actually receive PEFT updates. Even under this degree of sparsity, TS-PEFT is able to match or outperform the original baselines in commonsense reasoning, suggesting that standard PEFT exhibits substantial redundancy along the token dimension, and that threshold-based token-level selective updating provides a more efficient use of PEFT capacity.
@@
Visual Instruction Tuning. For visual instruction tuning, we follow the training setup of liu2024dora, configuring LoRA and DoRA with rank 128 and setting the initial and target ranks of AdaLoRA to 256 and 128. Table 2 shows the results. Across all benchmarks, TS-LoRA, TS-DoRA, and TS-AdaLoRA yield consistent improvements of about 0.3–0.4 points, while operating under 50–65% token sparsity. This demonstrates that token-level redundancy is also present in multimodal generative tasks, and selective token updates remain effective, consistent with our observations in CSR.
@@
Although TS-PEFT achieves a theoretical computation reduction of approximately 40%-60% in FLOPs within the adapter modules, current deep learning accelerators and libraries remain heavily optimized for dense matrix operations. Consequently, the imposed token-level sparsity currently serves as a theoretical computational bound and a probe for quantifying redundancy, rather than an immediate mechanism for wall-clock speedup. Nevertheless, with the emerging support for sparse tensor cores in next-generation hardware, TS-PEFT establishes a blueprint for translating this identified redundancy into tangible physical efficiency.
@@ 4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
Sparsity Patterns Reflect Architecture Bias. In LLaMA3.1-8B, there are 32 layers in total. We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.3). Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively. In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, showing a more “polarized” pattern. It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction. We hypothesize that this structural difference leads to the stronger selectivity observed in TS-DoRA at the module level, naturally raising the question: are the modules with lower sparsity more critical for downstream tasks?
@@
Redundant Updates Introduce Optimization Noise. To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conducted a cumulative selection analysis. We expanded the set of target modules in ascending order of sparsity (i.e., adding modules from “critical” to “redundant”). As illustrated in Figure 2, the conventional wisdom that ”more trainable parameters imply better performance” does not hold. For DoRA, performance peaks when only 50% of the modules are fine-tuned, while LoRA reaches its optimum at 80%. Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings
Crucially, we observe distinct “performance cliffs” when redundant modules are forced to update. For instance, DoRA suffers a sharp performance drop of 6.6% when the target proportion increases from 50% to 60%. Similarly, LoRA experiences a 1.3% degradation when expanding from 90% to 100%. This phenomenon provides compelling evidence that updating modules characterized by high sparsity injects negative interference rather than meaningful adaptation. In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism which automatically limits the capacity assigned to redundant positions.
@@ 5 Conclusion
This paper revisits parameter-efficient fine-tuning from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers. We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines. Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a theoretical basis for future hardware-software co-design targeting efficient sparse fine-tuning and inference.
@@ A.1 LLM Usage Statement
A.2 The training loss progression with and without the Adaptive Moment Estimation
Figure 3 illustrates the training dynamics of a single TS-PEFT layer (the v_proj in layers.20.self_attn of LLaMA3.1-8B) with and without Adam-style momentum on the threshold $\tau$. In both subfigures, the blue curve shows the training loss, while the red dashed curve tracks the corresponding threshold value over training steps (right-hand axis). When momentum is enabled (Figure 3(a)), the threshold increases smoothly and gradually saturates around a stable value, while the loss decreases rapidly at the beginning and then oscillates mildly around a low level. This indicates that $\tau$ can adapt to the task in a stable manner without causing large fluctuations in the optimization trajectory.
@@ A.2 The training loss progression with and without the Adaptive Moment Estimation
In contrast, when we remove the adaptive moment estimation and update $\tau$ using plain gradient descent (Figure 3(b)), the threshold trajectory becomes highly oscillatory and eventually drifts to a much larger range (above 0.6). The loss curve also exhibits large, persistent oscillations and fails to converge to a similarly low level as in Figure 3(a). This unstable behavior at the layer level is consistent with the degradation observed on downstream CSR benchmarks such as BoolQ: the dev accuracy drops noticeably compared with the full TS-PEFT variant that uses Adam-style updates for $\tau$. These results provide direct empirical evidence that adaptive momentum estimation is crucial for stabilizing the learning of token-level thresholds in practice.
A.3 Module-wise average sparsity of LLaMA3.1-8B on the commonsense reasoning task
Table 5 reports the detailed module-wise sparsity of TS-LoRA, TS-DoRA, and TS-AdaLoRA on LLaMA3.1-8B for commonsense reasoning. For each transformer layer (indexed from 0 to 31), we measure the token-level sparsity of the PEFT updates in the query, key, and value projections (q_proj, k_proj, and v_proj), and then compute their average (“avg” column). The last row summarizes the mean sparsity over all layers for each method. A dash (“–”) indicates that the corresponding module is not assigned any rank (i.e., effectively pruned by the underlying PEFT scheme).
From the table, we observe that TS-LoRA and TS-AdaLoRA exhibit relatively smooth sparsity profiles across layers: their per-layer averages mostly stay within the 50%–70% range, and no single layer dominates the overall sparsity pattern. This matches the moderate standard deviations reported in the main text (9.8 for TS-LoRA and 11.4 for TS-AdaLoRA), suggesting that these two methods tend to distribute token-level updates more evenly over depth. In contrast, TS-DoRA shows a much more polarized pattern: several layers (e.g., layers 0–2, 5–7, 11, 17, 24, 29) have near-100% sparsity in one or more projections, while a few layers (such as 8, 9, 14, 22) maintain comparatively low sparsity. This leads to a much larger standard deviation (41.3), indicating that TS-DoRA concentrates effective updates into a smaller subset of layers. These fine-grained statistics support our observation in the main paper that the token-level sparsity learned by TS-DoRA can serve as a strong indicator of module importance and motivates the subsequent experiments that select only the low-sparsity (high-importance) modules for fine-tuning.
@@
A.4 Hyperparameters
Across all experiments, we tune the sparsity-related hyperparameters s and $\lambda$ separately from the other training configurations. For CSR and VIT, we fix the scaling factor at s = $4\times 10^{-5}$ and select $\lambda$ from the interval $10^{-7},10^{-3}]$ based on development performance. For NLU tasks, we instead set $s=1\times 10^{-4}$ and choose $\lambda$ from $10^{-8},10^{-3}]$. The remaining hyperparameters, including rank, scaling coefficient, dropout rate, optimizer, learning rate, scheduler, batch size, number of epochs, and target modules, are summarized in Table 6.

### v3

@@ 3.4 Adam-style Threshold Update and Overall Training Procedure
In fact, for the threshold $\tau$, its approximate gradient $g_{k}$ is highly noisy and varies in scale across training steps. The contributions of tokens change significantly between batches, and the magnitude of $g_{k}$ can fluctuate during optimization. Directly applying plain gradient descent to $\tau$ often leads to oscillations and can slow down or make the convergence unstable. To mitigate this, we adopt a rule similar to Adam to update, which combines momentum and adaptive learning rate Kingma et al. (2015).
@@
where $\beta_{1}$ controls the contribution of past gradients and $\beta_{2}$ governs the decay rate of squared gradients. $\hat{m}_{k}$ and $\hat{v}_{k}$ denote the bias-corrected first- and second-moment estimates used in the Adam update. Finally, the threshold is updated as:
@@
where $s$ is the gradient scaling factor, and $\epsilon$ is a minuscule constant to prevent division by zero. We empirically observe that the adaptive moment estimation (Eq.(16)-Eq.(20)) is critical for convergence. Ablation studies in Appendix A.4 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks.
@@
Combining the above components, the training process of TS-PEFT can be summarized as follows: In each training step, during the forward propagation phase, we perform gating based on each token’s relative update magnitude $r_{i}$ and the threshold $\tau$ of the corresponding layer, applying updates only at positions that satisfy the condition while keeping the base model output unchanged at all other positions. Then, standard backpropagation is used to update all PEFT parameters, while the base model parameters remain frozen throughout. On this basis, we compute $\mu_{i}$ to construct the approximate gradient $g_{k}$ for the threshold, and independently update the threshold $\tau$ of each layer using Adam-style first- and second-moment estimates. The full training procedure is presented in Algorithm 1.
@@ 4.1 Experimental Setup
We evaluate TS-PEFT in multiple representative scenarios: Natural Language Understanding (NLU), Commonsense Reasoning (CSR), Visual Instruction Tuning (VIT), and Natural Language Generation (NLG). Due to space limitations, we present the detailed results and analysis for NLG in Appendix A.3. The experimental setup covers the datasets, baseline models, and hyperparameter configurations to ensure reproducibility and comparability of the results.
@@
We select several representative benchmarks in the current PEFT and small-model fine-tuning literature. For CSR, we evaluate on LLaMA-3.1-8B Dubey et al. (2024) using PIQA, BoolQ, HellaSwag, WinoGrande, SIQA, OBQA, and ARC-e/ARC-c, covering typical language reasoning scenarios such as physical commonsense, boolean question answering, narrative completion, and scientific QA Liu et al. (2024c). For VIT, we evaluate LLaVA-1.5-7B Liu et al. (2023, 2024a) on ScienceQA, POPE, MMBench, GQA, VisWiz, VQAv2, and VQAT, which are generative language tasks with open-ended answers conditioned on images and textual instructions. For NLU, we build on DeBERTaV3-base He et al. (2021) and use the GLUE benchmark datasets Wang et al. (2018). The GLUE tasks follow the official evaluation protocol: CoLA uses the Matthews correlation coefficient, STS-B uses correlation coefficients, and all other tasks use accuracy. Additionally, for the NLG tasks (detailed in Appendix), we evaluate Mistral-7B-v0.1 Jiang et al. (2023) on the Vicuna Chiang et al. (2023), LIMA Zhou et al. (2023) and Dolly Conover et al. (2023) datasets.
@@
In our experiments, we select several mainstream PEFT methods as baselines, covering two major categories: low-rank update methods and adapter-based methods. Specifically, LoRA, DoRA, and AdaLoRA are used to examine whether TS-PEFT can consistently provide additional gains under different low-rank update strategies, while Parallel Adapter Hu et al. (2023) is included to evaluate the method’s scalability on adapter-style architectures. For each baseline, we keep its original implementation, hyperparameters, and rank configuration unchanged, and apply TS-PEFT on top of it. This ensures that the comparison is not affected by factors unrelated to the baseline itself and that the results accurately reflect the impact of token-level selective updates.
@@
In all experiments, we freeze the base model and update only the PEFT parameters and the threshold $\tau$. We use the AdamW optimizer, where PEFT is the default learning rate and the effective learning rate of $\tau$ is further modulated by the scaling factor $s$. For text tasks, we follow standard instruction-tuning configurations. For vision–language tasks, we adopt the original training pipeline of the corresponding models. Unless otherwise specified, we set $\alpha_{k}=1$, $\beta_{1}=0.9$, and $\beta_{2}=0.98$. Each main experiment is repeated at least three times with different random seeds, and we report the mean scores across runs. Additional hyperparameter settings are summarized in Appendix A.6.
@@ 4.2 Experimental Results
Commonsense Reasoning. Table 1 reports the results of various PEFT methods and their TS-PEFT variants on the CSR tasks. For a fair comparison, we set the rank of both LoRA and DoRA to 32, and configure AdaLoRA with an initial rank of 64 and a target rank of 32. Under the same rank settings, we observe that TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve performance comparable to or better than their respective baselines on most datasets. This indicates that selectively skipping updates on a subset of tokens does not weaken downstream performance. Instead, it can help alleviate the noise introduced by redundant updates.
The table also reports the sparsity rate of TS-PEFT, which is the proportion of positions in the input sequence that are not modified by the PEFT output, as defined in Eq.(21). Across different baselines and tasks, this proportion remains roughly within the 50%–70% range, meaning that only 30% to 50% of the tokens actually receive PEFT updates. Even under this degree of sparsity, TS-PEFT is able to match or outperform the original baselines in commonsense reasoning, suggesting that standard PEFT exhibits substantial redundancy along the token dimension, and that threshold-based token-level selective updating provides a more efficient use of PEFT capacity.
@@
Visual Instruction Tuning. For visual instruction tuning, we follow the training setup of Liu et al. (2024c), configuring LoRA and DoRA with rank 128 and setting the initial and target ranks of AdaLoRA to 256 and 128. Table 2 shows the results. In all benchmark tests, TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve consistent improvements of about 0.3–0.4 points, while operating under 50%–70% token sparsity. This demonstrates that token-level redundancy is also present in multimodal generative tasks, and selective token updates remain effective, consistent with our observations in CSR.
@@
TS-PEFT has achieved PEFT updates for only 30% - 70% of the tokens in multiple tasks, yet still maintain or even surpass the baseline performance. This clearly reveals the widespread token-level redundant updates in the current PEFT mode, and these updates also introduce noise that affects model performance. By quantifying this intrinsic redundancy, TS-PEFT establishes a clear blueprint for translating theoretical sparsity into substantial physical efficiency gains in the future.
@@ 4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
Sparsity Patterns Reflect Architecture Bias. There are 32 layers in LLaMA3.1-8B. We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.5). Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively. In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, which shows a more polarized pattern. It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction. We hypothesize that this structural difference leads to the stronger selectivity observed in TS-DoRA at the module level, naturally raising the question: Are the modules with lower sparsity more critical for downstream tasks?
@@
Redundant Updates Introduce Optimization Noise. To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conduct a cumulative selection analysis. We expand the set of target modules in ascending order of sparsity. Specifically, we add the modules from the key ones to the redundant ones sequentially. As illustrated in Figure 2, the conventional wisdom that “more trainable parameters imply better performance” does not hold. For DoRA, performance peaks when only 50% of the modules are fine-tuned, while LoRA reaches its optimum at 80%. Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings.
Crucially, we observe distinct performance cliffs when redundant modules are forced to update. For instance, DoRA suffers a sharp performance drop of 6.6% when the target proportion increases from 50% to 60%. Similarly, LoRA experiences a 1.3% degradation when expanding from 90% to 100%. This phenomenon provides compelling evidence that updating modules characterized by high sparsity injects negative interference rather than meaningful adaptation. In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism.
@@ 5 Conclusion
This paper revisits PEFT from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers. We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 30%–70% of token positions, yet consistently matches or surpasses strong baselines. Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a theoretical basis for future hardware-software co-design targeting efficient sparse fine-tuning and inference.
@@ A.1 LLM Usage Statement
A.2 Case Study
Figure 3: Visualization of token-level sparsity learned by TS-LoRA on a BoolQ sample.
To explicitly demonstrate the token selection of TS-PEFT, we visualize the learned token-level sparsity on a representative validation sample from the BoolQ dataset, as shown in Figure 3.
By observing the distribution of the sparsity patterns, we find that TS-PEFT exhibits a clear semantic filtering mechanism when processing BoolQ tasks, which is manifested in three aspects:
  • Suppression of High-Frequency Function Words and Syntactic Noise: The model automatically freezes high-frequency function words and stop words such as “is”, “that”, “the”, and “:”. These words already possess well-established syntactic functions from the pre-training phase, and forcibly fine-tuning them often introduces unnecessary optimization noise. TS-PEFT effectively preserves the linguistic knowledge of the pre-trained model by skipping these positions.
  • Focus on Critical Domain Entities: For entity words closely related to the core logic of the question, such as “ethanol”, “energy”, and “produces”, the model chooses to update their parameters. This indicates that TS-PEFT can accurately isolate information carriers and concentrate the optimization budget thereon.
  • Preservation of Generic Pre-trained Concepts: Interestingly, the model also tends to keep the core label words in the output format, such as “true”, “false”, frozen. This implies that the model can reuse the truth-value concepts established during pre-training without re-adapting them during fine-tuning.
In conclusion, these qualitative results strongly support the hypotheses presented in the main text. We show that TS-PEFT is capable of acting as a dynamic probe that can autonomously identify and filter out redundant parts in the token level, thereby reducing the noise introduced by fine-tuning in inference tasks.
A.3 Additional Results
To further evaluate the generalization capability of TS-PEFT in open-ended text generation scenarios, we extend our experiments to NLG tasks. Specifically, we perform fine-tuning on the Alpaca dataset using the Mistral-7B-v0.1 model and subsequently evaluate performance on three representative benchmarks: Vicuna, LIMA, and Dolly. To measure generation quality comprehensively and objectively, we employed DeepSeek-R1-0325 as an automated evaluator, scoring model outputs on a 10-point scale across four dimensions: accuracy, relevance, fluency, and completeness.
As shown in Table 5, variants of TS-PEFT consistently outperform their corresponding dense update baselines across all test sets. Specifically, TS-LoRA, TS-DoRA, and TS-AdaLoRA surpass their baselines by average scores of 0.11, 0.10, and 0.12, respectively, while maintaining a token-level sparsity of approximately 75%, that is, updating only about one-quarter of the tokens. This result not only demonstrates the robustness of TS-PEFT in generation tasks but also provides strong empirical support for the core hypothesis of this paper. That is, in the process of fine-tuning, updating a vast number of non-critical tokens is not only computationally redundant but may even introduce noise, whereas dynamically filtering and skipping these updates via TS-PEFT facilitates the generation of higher-quality and more logically coherent responses.
Hyperparams Vicuna LIMA Dolly Avg
LoRA $s=4\mathrm{e}{-5}$ $\lambda=1\mathrm{e}{-5}$ 6.78 7.13 8.00 7.30
TS-LoRA 6.86 7.20 8.18 7.41
Sparsity(%) 76.8 78.8 76.2 77.3
DoRA $s=4\mathrm{e}{-5}$ $\lambda=6\mathrm{e}{-5}$ 7.02 6.95 7.82 7.26
TS-DoRA 7.09 6.83 8.15 7.36
Sparsity(%) 75.8 76.4 73.2 75.1
AdaLoRA $s=4\mathrm{e}{-5}$ $\lambda=6\mathrm{e}{-5}$ 6.94 6.74 7.59 7.09
TS-AdaLoRA 6.98 6.76 7.89 7.21
Sparsity(%) 72.2 77.3 75.3 74.9
Table 5: Evaluation scores of Mistral-7B-v0.1 on NLG benchmarks; the last row in each block shows the sparsity of TS-PEFT.
A.4 The training loss progression with and without the Adaptive Moment Estimation
Figure 4 illustrates the training dynamics of a single TS-PEFT layer (the v_proj in layers.20.self_attn of LLaMA3.1-8B) with and without Adam-style momentum on the threshold $\tau$. In both subfigures, the blue curve shows the training loss, while the red dashed curve tracks the corresponding threshold value over training steps. As shown in Figure 4 (a), when momentum is enabled, the threshold increases smoothly and gradually saturates around a stable value, while the loss decreases rapidly at the beginning and then oscillates mildly around a low level. This indicates that $\tau$ can adapt to the task in a stable manner without causing large fluctuations in the optimization trajectory.
@@ A.4 The training loss progression with and without the Adaptive Moment Estimation
In contrast, when we remove the adaptive moment estimation and update $\tau$ using plain gradient descent, as shown in Figure 4 (b), the threshold trajectory becomes highly oscillatory and eventually drifts to a much larger range (above 0.6). The loss curve also exhibits large, persistent oscillations and fails to converge to a similarly low level as in Figure 4(a). This unstable behavior at the layer level is consistent with the degradation observed on downstream CSR benchmarks such as BoolQ. That is, the accuracy drops noticeably compared with the TS-PEFT that uses Adam-style updates for $\tau$. These results provide direct empirical evidence that adaptive momentum estimation is crucial for stabilizing the learning of token-level thresholds in practice.
A.5 Module-wise average sparsity of LLaMA3.1-8B on the commonsense reasoning task
Table 6 reports the detailed module-wise sparsity of TS-LoRA, TS-DoRA, and TS-AdaLoRA on LLaMA3.1-8B for CSR benchmarks. For each transformer layer, that is, indexed from 0 to 31, we measure the token-level sparsity of the PEFT updates in the query, key, and value projections (q_proj, k_proj, and v_proj), and then compute their average, the “avg” column in the figure. The last row summarizes the mean sparsity over all layers for each method.
From the table, we observe that TS-LoRA and TS-AdaLoRA exhibit relatively smooth sparsity profiles across layers. Their per-layer averages mostly stay within the 50%–70% range, and no single layer dominates the overall sparsity pattern. This matches the moderate standard deviations reported in the main text (9.8 for TS-LoRA and 11.4 for TS-AdaLoRA), suggesting that these two methods tend to distribute token-level updates more evenly over depth. In contrast, TS-DoRA shows a much more polarized pattern: several layers (e.g., layers 0–2, 5–7, 11, 17, 24, 29) have near-100% sparsity in one or more projections, while a few layers (such as 8, 9, 14, 22) maintain comparatively low sparsity. This resulted in a relatively large standard deviation of 41.3, indicating that TS-DoRA concentrates effective updates into a smaller subset of layers. These fine-grained statistics support our observation in the main paper that the token-level sparsity learned by TS-DoRA can serve as a strong indicator of module importance and motivates the subsequent experiments that select only the low-sparsity (high-importance) modules for fine-tuning.
@@
A.6 Hyperparameters